OpenStudy (anonymous):

What are the solutions of 3x2 + 4x = -5? a. the quantity negative 4 plus or minus i square root of 66 all over 6 b. the quantity negative 2 plus or minus 2i square root of 66 all over 3 c. the quantity negative 2 plus or minus i square root of 11 all over 3 d. the quantity negative 4 plus or minus i square root of 11 all over 6

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

for the general equation ax^2 +bx+c, \[x= \frac{ -b \pm \sqrt{b ^{2}-4*a*c} }{ 2*a }\]

OpenStudy (anonymous):

yah i got \[-4±\sqrt{4^2-4*(3)*(5)}\div6\]

OpenStudy (anonymous):

well over 6

OpenStudy (anonymous):

yeah you are doing good what will be next step?

OpenStudy (anonymous):

simlify right?

OpenStudy (anonymous):

yeah do the simplification

OpenStudy (anonymous):

\[x=\frac{ -2±\sqrt{16-60} }{3 }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@niksva

OpenStudy (anonymous):

\[\frac{ -4\pm \sqrt{16-60} }{ 6 }\]

OpenStudy (anonymous):

\[\frac{ -4\pm \sqrt{-44} }{ 6 }\] \[\frac{ -4\pm2\sqrt{11}i }{ 6 }\] \[\frac{ -2\pm \sqrt{11}i }{ 3 }\]

OpenStudy (anonymous):

ohhhhhh ok thank you so much man :)

OpenStudy (anonymous):

I got a 100% on my quiz bc you helped me @niksva :)))))))))))))